OpenStudy (loser66):

x^2y"+y=0 \(\alpha =0; \beta=1\) therefore, characteristic equation is \(r^2-r+1=0\) and \(r_{1,2} = \frac{1}{2}\pm\frac {\sqrt{3}}{2}i\) which leads to solution \[y=C_1x^{\frac{1}{2}}cos(\frac{\sqrt{3}}{2}lnx)+C_2x^{\frac{1}{2}sin(\frac{\sqrt{3}}{2}lnx)\] Then, I don't know how to do next. Please, help

OpenStudy (loser66):

\[r^2-r+1=0\\r=\frac{1}{2}\pm\frac{\sqrt{3}}{2}i\]

OpenStudy (loser66):

\[y=C_1x^{\frac{1}{2}}cos(\frac{\sqrt{3}}{2}lnx) +C_2 x^{\frac{1}2}sin(\frac{\sqrt{3}}{2}lnx)\]

OpenStudy (loser66):

then, I am stuck, @oldrin.bataku help me, please

OpenStudy (experimentx):

try putting y(x) into your DE and see if it is solution or not

OpenStudy (loser66):

I shortcut a step which tests whether x =0 is regular singular point or not. It passes that test, that mean that y is solution.

OpenStudy (loser66):

\[lim_{x\rightarrow 0}x^2\frac{1}{x^2} =1\]

OpenStudy (experimentx):

whats the problem?

OpenStudy (loser66):

I need a solution under the form of series \(\sum\)

OpenStudy (experimentx):

the series method is a mess, this is just Euler-Cauchy equation, that it is best dealt with. if you want final answer in series then just expand the final answer.

OpenStudy (loser66):

how?

OpenStudy (experimentx):

assume the solution be of the form y=x^n, n(n-1) + 1 = 0 , let n1 and n2 be the roots of this quadratic equation your solution should be of the form y(x) = C1x^(n1) + C2 x^(n2)

OpenStudy (experimentx):

http://en.wikipedia.org/wiki/Frobenius_method since 1/x^2 is not analytic at x=0, no power series solution exist. \[ \lim_{x\to 0} x^{1/2} \sin( k \log(x)) = 0\] http://www.wolframalpha.com/input/?i=limit+x-%3E0+x%5E%281%2F2%29+sin%28log%28x%29%29 y[x] = 0 satisfield eqn for any x, if y[x] is singular at x=0, then probably that's the case of removable singularity.